OpenStudy (anonymous):

Given triangle DOG, where angle D=64 degrees, angle O= 38 degrees, length of DO = 8 inches. I need to solve the problem completely and find the area of triangle DOG. Any help ?

OpenStudy (anonymous):

Sorryi was terrbile in geometry.

OpenStudy (anonymous):



OpenStudy (anonymous):

G you find first because angles have to add up to 180

OpenStudy (anonymous):

So G would be 78?

OpenStudy (anonymous):

D=64 O=38 G= 78

OpenStudy (anonymous):

Then I have to use sine right ?

OpenStudy (anonymous):

you find the other sides via law of sines

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but to find the other sides, would that be one expression or two? Can I find each side separately versus finding them at the same time?

OpenStudy (anonymous):

so for example \[\frac{8}{\sin(78)}=\frac{d}{\sin(64)}\] and threfore \[d=\frac{8\sin(64)}{\sin(78)}\]

OpenStudy (anonymous):

where d is side opposite D you find them one at a time

OpenStudy (anonymous):

now a calculator

OpenStudy (anonymous):

i get 7.35

OpenStudy (anonymous):

now you find o same method.

OpenStudy (anonymous):

okay i got that. what about the other side?

OpenStudy (anonymous):

i think i know what you mean....let me try it.

OpenStudy (anonymous):

same method. what i would call o and you could call DG use law of sines again. gotta run but i bet you can get it. same as before

OpenStudy (anonymous):

thanks !